# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially-Private Stochastic Saddle-Point Problems

### v1

@@ 4.1 Private Stochastic Mirror Descent with Sampling from Vertices
Let $(\tilde{x},\tilde{y})$ be the output of Algorithm 1 and define $(\bar{x},\bar{y})$ as in Lemma 4.1. Then,
$\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq\frac{4L_{1}}{T}+\frac{2L_{1}\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{T}}.$
With these results, we are ready to prove Theorem 4.1.
Proof of Theorem 4.1.
Denote $(\tilde{x},\tilde{y})=\frac{1}{T}\sum_{t=1}^{T}(\tilde{x}^{t},\tilde{y}^{t})$ the private output of Algorithm 1, and $(\bar{x},\bar{y})=\frac{1}{T}\sum_{t=1}^{T}(x^{t},y^{t})$ its ‘non-private’ average. Using Lemma 4.1 and the fact that $\|g^{t}\|_{\infty}^{2}+\|\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\|_{\infty}^{2}\leq 5L_{0}^{2}$ by $L_{0}$-Lipschitzness of the objective we obtain
$\operatorname{Gap}(\bar{x},\bar{y})\leq\frac{2(\log(d_{x})+\log(d_{y}))}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle.$
@@
Since $w^{t},v^{t}$ are deterministic when conditioning on the randomness up to iteration $t$, taking expectations on both sides of the previous inequality leads to
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]$ $\lesssim\frac{2(\log(d_{x})+\log(d_{y}))}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}.$ (1)
@@
Since we are assuming second order smoothness, then $|\mathbb{E}[\nabla_{x,j}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x,j}\mid\mathcal{F}_{t}]|\leq\frac{2L_{2}}{K}$ by Corollary 3.1. Similarly, for any $i\in[d_{y}]$ we have $|\mathbb{E}[-\nabla_{y,i}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y,i}\mid\mathcal{F}_{t}]|\leq 2L_{2}/K$. We conclude that $\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}\leq 2L_{2}/K$. In addition, from Lemma 4.2 it follows that $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq 8L_{1}\sqrt{\log(d_{x})+\log(d_{y})}/\sqrt{T}$. These bounds lead to
$\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]\lesssim L_{1}\frac{\sqrt{\log(d_{x})+\log(d_{y})}}{\sqrt{T}}+\frac{(\log(d_{x})+\log(d_{y}))}{\tau T}+\tau L_{0}^{2}+\frac{L_{2}}{K}.$ (2)
We obtain the first convergence rate by optimizing this bound over $T,\tau,L$, taking into account the constraint over $\tau$ that Proposition 4.1 requires for the privacy. ∎
@@
4.2 Private Stochastic Mirror Descent with Bias-Reduced Sampling from Vertices
@@
The limitation that prevents Algorithm 1 from attaining optimal rates is the bias of the stochastic gradient oracles. This is a consequence of sampling from vertices, and can only be mitigated by increasing the sampling rate. However, this eats up the privacy budget. In this section, we propose circumventing this limitation by combining our gradient oracles with a bias-reduction method.
Bias-Reduced Gradient Oracle.
@@
We say a random variable $N$ is distributed as a truncated geometric with parameters $(p,M)$, denoted by $\mbox{TG}(p,M)$, if $\mathbb{P}[N=k]=p^{k}\mathbbm{1}_{(k\in\{0,1,...,M\})}/C_{M}$, where the normalization constant is $p\leq C_{M}\leq 1$. To construct a bias-reduced estimator of the gradient at $(x,y)$, we start by sampling $N\sim\mbox{TG}(0.5,M)$. Then, we construct a batch $B$ of $\max\{1,2^{N^{t}}/\alpha\}$ fresh samples, where $\alpha$ is a parameter to be choosen later. This batch $B$ combined with the gradient estimates (see Algorithm 2 and note similar algorithms in Blanchet and Glynn (2015)) reduces the bias while preserving privacy and accuracy (see Thm. 4.3). To do so, we control the bias and second moment of the gradient estimator. The proof of the next result is deferred to Appendix E.2.
@@
[…]
  $(\bar{x}_{+},\bar{y}_{+})=\frac{1}{2^{N+1}}\sum_{i\in[2^{N+1}]}(\hat{x}^{i},\hat{y}^{i})$ ; $(\bar{x}_{-},\bar{y}_{-})=\frac{1}{2^{N}}\sum_{i\in[2^{N}]}(\hat{x}^{i},\hat{y}^{i})$ ; $(x_{0},y_{0})=(\hat{x}^{1},\hat{y}^{1})$
  $g_{x}=C_{M}2^{N}(\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B))+\nabla_{x}f(x_{0},y_{0};B)$
  $g_{y}=C_{M}2^{N}(\nabla_{y}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{y}f(\bar{x}_{-},\bar{y}_{-};B))+\nabla_{y}f(x_{0},y_{0};B)$
@@
[…]
$\|\mathbb{E}[\Phi_{\mathcal{D}}(x,y)-(g_{x},-g_{y})]\|_{\infty}$ $\leq\frac{2L_{2}}{2^{M}},$
[…]
@@
Consequently, the bias scales inversely proportional to $2^{M}$, where $M$ is largest possible realization of $N$. Hence, the expected bias is the same as if we were sampling approximately $2^{M}$ times while we actually sample only $2^{N}$ times. This idea leads to a more favorable privacy-utility tradeoff. Another consequence is the order of the second moment, $M(\log(d_{x})+\log(d_{y}))$. Here $M$ will be choosen as a logarithmic quantity. Hence, the method allows us to improve the bias at a only a logarithmic cost in the second moment bound. Note, however, a disadvantage of this gradient oracle is we can only control the second moment for second-order-smooth objective functions.
Bias-reduced Algorithm, Algorithm 3.
@@
We propose an algorithm which is essentially Algorithm 1 combined with the bias-reduced gradient oracle. The main challenges encountered are the oracles becoming heavy-tailed (moments are exponentially larger than their expectation), and due to a randomization of batch sizes, it is difficult to control the privacy budget. To handle the heavy-tailed nature of our estimators, we resort to efficiency estimates from Markov’s inequality which only certify success with constant probability. On the other hand, the privacy budget analysis uses the Fully Adaptive Composition Theorem of DP (Whitehouse et al., 2023), which allows us in particular to set up the mechanisms and their privacy parameters in a random (but predictable) way.
@@
Algorithm 3 Private stochastic mirror descent with bias-reduced sampling from vertices

### v2

@@ 3.1.2 Convergence Analysis of Algorithm 1
[…]
Proof 3.12 (Proof of Lemma 3.11).
Let $F_{\lambda}(x):=\max_{y\in\Delta_{y}}[F_{\mathcal{D}}(x,y)+\lambda\phi(y)]$ where $\phi$ is the negative entropy. Since $F_{\mathcal{D}}(x,y)$ is $L_{1}$-smooth in $(x,y)$ and concave in $y$, and $\lambda\phi(y)$ is $\lambda$-strongly convex in $y$, by Proposition 1 in [generalized_fenchel-young_losses], $F_{\lambda}$ is $\big(L_{1}+\frac{L_{1}^{2}}{\lambda}\big)$-smooth w.r.t $\|\cdot\|_{1}$. Hence by Corollary 2.3 applied to $F_{\lambda}$, we have $\mathbb{E}[F_{\lambda}(\tilde{x})-F_{\lambda}(\bar{x})]\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}$. In addition, it is easy to see that $\max_{v\in\Delta_{y}}F_{\mathcal{D}}(x,v)\leq F_{\lambda}(x)\leq\lambda\log(d_{y})+\max_{v\in\Delta_{y}}F_{\mathcal{D}}(x,v)$ for all $x\in\Delta_{x}$, since $0\leq\phi(y)\leq\log(d_{y})$. Therefore,
$\mathbb{E}\big[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)\big]=\mathbb{E}\big[\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\tilde{x},v)-F_{\lambda}(\tilde{x})\big]+\mathbb{E}[F_{\lambda}(\tilde{x})-F_{\lambda}(\bar{x})]\\ +\mathbb{E}\big[F_{\lambda}(\bar{x})-\max_{v\in\Delta_{y}}F_{\mathcal{D}}(\bar{x},v)\big]\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{y}).$
Analogously, $\mathbb{E}\big[\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\bar{y})-\min_{w\in\Delta_{x}}F_{\mathcal{D}}(w,\tilde{y})\big]\leq\frac{2L_{1}}{T}+\frac{2L_{1}^{2}}{\lambda T}+\lambda\log(d_{x})$. Using both inequalities, we get
[…]
$=\frac{4L_{1}}{T}+\frac{4L_{1}^{2}}{\lambda T}+\lambda\ell+\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})].$
Choosing $\lambda=L_{1}\frac{2}{\sqrt{T\ell}}$ gives the bound.
With these results, we are ready to prove the convergence guarantees.
Proof 3.13 (Convergence of Algorithm 1).
@@
Following the notation in Algorithm 1, denote $(\tilde{x},\tilde{y})=\frac{1}{T}\sum_{t=1}^{T}(\tilde{x}^{t},\tilde{y}^{t})$ and $(\bar{x},\bar{y})=\frac{1}{T}\sum_{t=1}^{T}(x^{t},y^{t})$. Using Lemma 3.8 and the fact that $\|g^{t}\|_{\infty}^{2}+\|\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\|_{\infty}^{2}\leq 5L_{0}^{2}$ by $L_{0}$-Lipschitzness of the objective, we obtain $\operatorname{Gap}(\bar{x},\bar{y})\leq\frac{2\ell}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\langle\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t},(x^{t},y^{t})-(w^{t},v^{t})\rangle$. Since $w^{t},v^{t}$ are deterministic when conditioning on the randomness up to iteration $t$, taking expectations on both sides of the previous inequality leads to
(8) $\mathbb{E}[\operatorname{Gap}(\bar{x},\bar{y})]$ $\lesssim\frac{2\ell}{\tau T}+5\tau L_{0}^{2}+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}.$
[…]
@@
Now, we prove Equation (6). Note that $|\mathbb{E}[\nabla_{x,j}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{x,j}\mid\mathcal{F}_{t}]|\leq\frac{2L_{2}}{K}$ by Corollary 2.4 and the second-order-smoothness assumption. Similarly, we have $|\mathbb{E}[-\nabla_{y,i}F_{\mathcal{D}}(x^{t},y^{t})-g^{t}_{y,i}\mid\mathcal{F}_{t}]|\leq\frac{2L_{2}}{K}$ for all $i\in[d_{y}]$. We conclude that $\|\mathbb{E}[\Phi_{\mathcal{D}}(x^{t},y^{t})-g^{t}\mid\mathcal{F}_{t}]\|_{\infty}\leq\frac{2L_{2}}{K}$. Moreover, by Lemma 3.11, $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})-\operatorname{Gap}(\bar{x},\bar{y})]\leq 8L_{1}\sqrt{\ell}/\sqrt{T}$. Hence $\mathbb{E}[\operatorname{Gap}(\tilde{x},\tilde{y})]\lesssim L_{1}\frac{\sqrt{\ell}}{\sqrt{T}}+\frac{\ell}{\tau T}+\tau L_{0}^{2}+\frac{L_{2}}{K}$. Setting $T\eqsim\min\Big\{n,\frac{(n\varepsilon)^{4/5}}{\ell^{1/5}\log(1/\delta)^{2/5}}\Big\}$, $\tau\eqsim\min\Big\{\frac{1}{L_{0}}\sqrt{\frac{\ell}{T}},\frac{n\varepsilon}{L_{0}T\sqrt{TK\log(1/\delta)}}\Big\}$ and $K\eqsim\sqrt{\frac{T}{\ell}}$ proves (6).
Finally, we prove Equation (7). Since $L_{2}=0$, we can safely set $K=1$. Optimizing over $T,\tau$ gives the bound.
@@
3.2 A constant probability duality gap bound via Private Stochastic Mirror Descent with Bias-Reduced Sampling from Vertices
@@
The limitation that prevents Algorithm 1 from attaining optimal rates is the bias of the stochastic gradient oracles. This is a consequence of sampling from vertices, and can only be mitigated by increasing the sampling rate. However, this eats up the privacy budget. In this section, we propose circumventing this limitation by combining our gradient oracles with a bias-reduction method; see Algorithm 3. We first show how that this procedure leads to improved rates, but only with constant probability. In Section 3.3 we show how to obtain a high probability bound of roughly the same order.
Theorem 3.14 (Guarantees for Algorithm 3).
[…]
$\mathbb{P}\left[\operatorname{Gap}(\tilde{x},\tilde{y})\lesssim\log(n)\bigg[\sqrt{C}\sqrt{\frac{\ell}{n}}+\sqrt{L_{0}}C^{1/4}\left(\frac{\ell^{3/2}\sqrt{\log(1/\delta)}}{n\varepsilon}\right)^{1/2}\bigg]\right]\geq 0.99.$
Up to poly-logarithmic factors in $n,d_{x},d_{y}$, the rate above is indeed optimal. Before proving this result, we will explain Algorithm 3 in more detail.
Bias-Reduced Gradient Oracle
@@
We say a random variable $N$ is distributed as a truncated geometric with parameters $(p,M)$ if $\mathbb{P}[N=k]=p^{k}\mathbbm{1}_{(k\in\{0,1,...,M\})}/C_{M}$, where the normalizing constant is $p\leq C_{M}\leq 1$. We denote it by $\mbox{TG}(p,M)$. To construct a bias-reduced estimator of the gradient at $(x,y)$, we start by sampling $N\sim\mbox{TG}(0.5,M)$. Then, we construct a batch $B$ of $\max\{1,2^{N^{t}}/\alpha\}$ fresh samples, where $\alpha$ is a parameter to be choosen later. This batch $B$ combined with the gradient estimates (see Algorithm 2 and note similar algorithms in [Blanchet:2015]) reduces the bias while preserving privacy and accuracy (see Thm. 3.14). To do so, we control the bias and second moment of the gradient estimator.
@@
 Sample $(\hat{x}^{i},\hat{y}^{i})_{i\in[2^{N+1}]}\overset{\text{iid}}{\sim}P_{x}\times P_{y}$
$(\bar{x}_{+},\bar{y}_{+})=\frac{1}{2^{N+1}}\sum_{i\in[2^{N+1}]}(\hat{x}^{i},\hat{y}^{i})$ ; $(\bar{x}_{-},\bar{y}_{-})=\frac{1}{2^{N}}\sum_{i\in[2^{N}]}(\hat{x}^{i},\hat{y}^{i})$
$g_{x}=C_{M}2^{N}(\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B))+\nabla_{x}f(\hat{x}^{1},\hat{y}^{1};B)$
$g_{y}=-C_{M}2^{N}(\nabla_{y}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{y}f(\bar{x}_{-},\bar{y}_{-};B))-\nabla_{y}f(\hat{x}^{1},\hat{y}^{1};B)$
@@
Suppose the objective of the SSP problem in the $\ell_{1}$ setting is second-order-smooth. Let $x\in\Delta_{x},y\in\Delta_{y}$, $N\sim\mbox{TG}(0.5,M)$ and $B\sim\mathcal{D}^{N}$. Let $(g_{x},g_{y})=\text{BiasReducedGradient}(x,y,N,B)$. Then,
$\|\mathbb{E}[\Phi_{\mathcal{D}}(x,y)-(g_{x},g_{y})]\|_{\infty}$ $\leq\frac{2L_{2}}{2^{M}},$
$\text{and}\quad\mathbb{E}[\|(g_{x},g_{y})\|_{\infty}^{2}]$ $\lesssim L_{0}^{2}+L_{2}^{2}+L_{1}^{2}M[\log(d_{x})+\log(d_{y})].$
@@
Consequently, the bias scales inversely proportional to $2^{M}$, where $M$ is largest possible realization of $N$. Hence the expected bias is roughly same as if we were sampling $2^{M}$ times while we actually sample only $2^{N}$ times. This leads to a more favorable privacy-utility tradeoff. Another consequence is the order of the second moment, $M\ell$. Here $M$ will be choosen as a logarithmic quantity (i.e., $M\lesssim\log(\min\{n\varepsilon,n\})$). Hence, the method allows us to improve the bias at a only a logarithmic cost in the second moment bound. Note, however, a disadvantage of this gradient oracle is we can only control its second moment for second-order-smooth objective functions.
Proof 3.16 (Proof of Lemma 3.15).
Following the notation in Algorithm 2, we have
$\mathbb{E}[g_{x}]=\textstyle\mathbb{E}\left[\sum_{k=0}^{M}\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)+\frac{2^{-k}}{C_{M}}\nabla_{x}f(\hat{x}^{1},\hat{y}^{1};B)\mid N=k\right]$
$=\textstyle\mathbb{E}\left[\sum_{k=0}^{M}\big(\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+},\bar{y}_{+})-\nabla_{x}F_{\mathcal{D}}(\bar{x}_{-},\bar{y}_{-})\big)+\frac{2^{-k}}{C_{M}}\nabla_{x}F_{\mathcal{D}}(\hat{x}^{1},\hat{y}^{1})\mid N=k\right]$
$=\textstyle\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+},\bar{y}_{+})\mid N=M\right]-\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(x_{-},y_{-})\mid N=0\right]+\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\hat{x}^{1},\hat{y}^{1})\right]$
$=\textstyle\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+},\bar{y}_{+})\mid N=M\right],$
where the first equality follows from the law of total expectation, the second from the fact that $(\bar{x}_{+},\bar{y}_{+})$$(\bar{x}_{-},\bar{y}_{-})$, $(\hat{x}^{1},\hat{y}^{1})$ and $B$ are conditionally independent given $N$, the third from $\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{+},\bar{y}_{+})\mid N=k\right]=\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\bar{x}_{-},\bar{y}_{-})\mid N=k+1\right]$ and the last one from $\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(x_{-},y_{-})\mid N=0\right]=\mathbb{E}\left[\nabla_{x}F_{\mathcal{D}}(\hat{x}^{1},\hat{y}^{1})\right]$. Finally, we get $\|\mathbb{E}[g_{x}]-\nabla_{x}F_{\mathcal{D}}(x,y)\|_{\infty}=\max_{j\in[d]}|\mathbb{E}[g_{x,j}]-\nabla_{x,j}F_{\mathcal{D}}(x,y)|\leq\frac{2L_{2}}{2^{M}}$ by Corollary 2.3. Clearly, an analogous bound holds for the gradient with respect to $y$. Writing $\|\mathbb{E}[(g_{x},g_{y})]-\Phi_{\mathcal{D}}(x,y)\|_{\infty}=\max\{\|\mathbb{E}[g_{x}]-\nabla_{x}F_{\mathcal{D}}(x,y)\|_{\infty},\|\mathbb{E}[g_{y}]-\nabla_{y}F_{\mathcal{D}}(x,y)\|_{\infty}\}$ gives the bias bound.
We now proceed with the second moment bound. Then, $\mathbb{E}\|g_{x}\|_{\infty}^{2}$ is equal to
$\textstyle\sum_{k=0}^{M}\mathbb{E}\|C_{M}2^{k}[\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B_{k})-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)]+\nabla_{x}f(\hat{x}^{1},\hat{y}^{1};B)\mid N=k\|_{\infty}^{2}\frac{2^{-k}}{C_{M}}$
$\textstyle\leq 2\left[\sum_{k=0}^{M}\!C_{M}2^{k}\mathbb{E}\|\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)\!-\!\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)\mid N=k\|_{\infty}^{2}\!+\!\frac{2^{-k}}{C_{M}}L_{0}^{2}\right]$
$\textstyle\leq 2L_{0}^{2}+4C_{M}\sum_{k=0}^{M}2^{k}\mathbb{E}\|\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(x,y;B)\mid N=k\|_{\infty}^{2}$
$\textstyle\quad+4C_{M}\sum_{k=0}^{M}\mathbb{E}2^{k}\|\nabla_{x}f(x,y;B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)\mid N=k\|_{\infty}^{2}.$
By Corollary 2.12, $\mathbb{E}\|\nabla_{x}f(\bar{x}_{+},\bar{y}_{+};B)-\nabla_{x}f(x,y;B)\mid N=k\|_{\infty}^{2}\lesssim\frac{L_{2}^{2}}{2^{2(k+1)}}+\frac{\log(d_{x})L_{1}^{2}}{2^{k+1}}$, and similarly $\mathbb{E}\|\nabla_{x}f(x,y;B)-\nabla_{x}f(\bar{x}_{-},\bar{y}_{-};B)\mid N=k\|_{\infty}^{2}\lesssim\frac{L_{2}^{2}}{2^{2k}}+\frac{\log(d_{x})L_{1}^{2}}{2^{k}}$. Hence, $\mathbb{E}\|g_{x}\|_{\infty}^{2}\lesssim L_{0}^{2}+L_{2}^{2}+M\log(d_{x})L_{1}^{2}$. Analogously, $\mathbb{E}\|g_{y}\|_{\infty}^{2}\lesssim L_{0}^{2}+L_{2}^{2}+M\log(d_{y})L_{1}^{2}$.
Bias-reduced Algorithm
@@
We propose Algorithm 3, which is essentially Algorithm 1 combined with the bias-reduced gradient oracle. The main challenges encountered are the oracles becoming heavy-tailed (moments are exponentially larger than their expectation), and, due to a randomization of batch sizes, it is difficult to control the privacy budget. To handle the heavy-tailed nature of our estimators, we resort to efficiency estimates from Markov’s inequality which only certify success with constant probability. For handling the privacy budget challenge, we use the Fully Adaptive Composition Theorem of DP [Whitehouse:2023], which allows us to set up the mechanisms and their privacy parameters in a random (but predictable) way.
@@
Algorithm 3 Private stochastic mirror descent with bias-reduced vertex sampling